Problem 5S-14

The director of social services of a county has learned that the state has mandated additional information requirements. This will place an additional burden on the agency. The director has identified three acceptable alternatives to handle the increased workload. One alternative is to reassign present staff members, the second is to hire and train two new workers, and the third is to redesign current practice so that workers can readily collect the information with little additional effort. An unknown factor is the caseload for the coming year when the new data will be collected on a trial basis. The estimated costs for various options and caseloads are shown in the following table:

CASELOAD
Moderate High Very High
Reassign staff $ 60 * 50 70
New staff 80 80 80
Redesign collection 45 55 65

*Cost in $ thousands.


After contemplating the caseload question, the director of social services has decided that reasonable caseload probabilities are 0.20 for moderate, 0.30 for high, and 0.50 for very high.

a. Which alternative will yield the minimum expected cost?

  • Reassign

  • Redesign

  • or

  • New Staff



c. Determine the expected value of perfect information using an opportunity loss table. (Round your answers to the nearest whole number. Leave no cells blank - be certain to enter "0" wherever required.)

Opportunity loss table
Moderate($) High($) Very High($) EOL($)
Reassign
New Staff
Redesign
(0.20) (0.30) (0.50)


Expected value of perfect information           $ ___________?
Answer #1

(a)

The expected cost of 'Reassign' = 60*0.2 + 50*0.3 + 70*0.5 = $62 thousand

The expected cost of 'Redesign' = 80*0.2 + 80*0.3 + 80*0.5 = $80 thousand

The expected cost of 'New staff' = 45*0.2 + 55*0.3 + 65*0.5 = $58 thousand (Min)

So,

The minimum expected cost is yielded by the 'New staff' alternative.

(b)

Moderate High Very high EOL
Reassign 15,000 0 5,000 5,500
New staff 35,000 30,000 15,000 23,500
Redesign 0 5,000 0 1,500
0.2 0.3 0.5

Expected value of perfect information (EVPI) = Min (EOL) = $1,500